# semicircular, free Poisson and R-transform in high order

### v1

@@ 1 Introduction
This work follows [12] and we refer to that previous study for a more detailed introduction to freeness for tensors. In this paper, we further develop and expand upon this notion of tensorial freeness.
@@
In many respects, random tensors generalize random matrices to higher orders. This field is currently a vibrant area of research, driven by numerous applications in physics and computer science [2, 23, 10, 4, 22]. Hence, in this paper, we aim to establish analogous results in the framework of tensorial freeness. The principal contributions of this work are the following :
@@
where $\gamma$ is the number of connected components of $\mathfrak{b}$. The distribution of a collection of tensors $\mathcal{A}=\{A_{1},\ldots,A_{n}\}$ of possibly different order, $A_{k}\in\mathcal{E}^{N}_{p_{k}}$ is the collection of all trace maps $\mathfrak{b}((T_{v})_{v\in V})$ with $T_{v}\in\mathcal{A}$ and $\mathfrak{b}$ with compatible degrees. In particular, if $T\in\mathcal{E}^{N}_{p}$, the distribution of $T$ is the collection of trace invariants for $p$-regular trace maps, that is all vertices of the map belong to exactly $p$-edges.
We introduce now informally three types of $p$-regular trace maps useful for the rest of the introduction. A melon is a map with two vertices and $p$ edges between them. If $p$ is even, a bouquet is a map with one vertex and $p/2$ self-loops, and a multicycle is a map with $n\geq 1$ vertices $p/2$ edges between the $i$-th and the $i+1$-th ones for all $1\leq i\leq n$ ($n+1=1$). For $p$ odd, an odd multicycle is a map with $2n$ vertices and $(p+1)/2$ edges between the $12i-$ 1-th and the $2i$-th ones, $(p-1)/2$ edges between the $2i$-th and the $12i+$ 1-th ones, for all $1\leq i\leq n$ ($12n+1=$ 1).
@@
The notion of freeness detailed in [12] arises from a poset (partially ordered set) on maps, called by analogy the non-crossing poset. Loosely speaking, we say that a map $\mathfrak{b}^{\prime}$ is smaller than a map $\mathfrak{b}$ if they differ by a switch, that is exchanging the ending point of two edges in $\mathfrak{b}$, and $\mathfrak{b}^{\prime}$ has one more connected component. By transitivity, this gives a poset on the set of trace maps with a given number of vertices and a given sequence of degrees. The formal definitions are recall in Section 2 and an example is given in Figure 2. By Moebius inversion, it is possible to define free cumulants associated to this poset characterized by the implicit relation
@@ 1.1 Main results
For a tensor $T$ of order $p$, and $n\geq 1$, we define $m_{n}(T)$, the $n$-th moment of $T$, as the sum over all the $p$-regular connected rooted trace maps with $n$ vertices of the trace invariants where all vertices are decorated by $T$. Similarly, $\kappa_{n}(T)$ the $n$-th free cumulant of $T$ is the sum over all the $p$-regular connected trace maps with $n$ vertices of free cumulant of the trace map where all vertices are decorated by $T$. We set $m_{0}(T)=\kappa_{0}(T)=1$. We then define the corresponding formal power series in $\mathbb{C}{{\left[\left[z\right]\right]}}$ :
@@
We have as functionnal relation in $\mathbb{C}{{\left[\left[z\right]\right]}}$,
@@
We have not yet succeeded to give a direct argument of the existence of this measure by showing the positivity of these moments, nevertheless the following results of convergence can always be given in terms of convergence of the moment and vanishing of the variance instead of in terms of weak convergence of the measure.
If $a$ is a distribution on the $p$-regular combinatorial maps (that is $a:\mathcal{M}_{0}\mapsto\mathbb{R}$ with $a(\emptyset)=1$), we can speak of its moments $m_{n}(a)$ as the sum of $a(\mathfrak{b})$ over all the $p$-regular connected maps with $n$ vertices, and we denote $\mu_{a}$ as the measure on $\mathbb{R}$ having $(m_{n}(a))_{n\geq 0}$ as moments, if it exists.
@@
    if $p$ is even,
    $b_{p,t}(\kappa_{\mathfrak{b}})=\frac{t}{[(p/2)!]^{n}}\mbox{ if }\mathfrak{b}\mbox{ is a multicycle of length }n\mbox{ and }0\mbox{ otherwise,}$,
    if $p$ is odd,
    $b_{p,t}(\kappa_{\mathfrak{b}})=\frac{t}{[(p+1/2)!(p-1/2)!]^{n}}\mbox{ if }\mathfrak{b}\mbox{ is an odd multicycle of length }2n\mbox{ and }0\mbox{ otherwise.}$.
@@
$\kappa_{2}(\mu_{p})=1\mbox{ if $n=0$ or $2$}\quad\quad\mathrm{and}\quad\quad\kappa_{n}(\mu_{p})=0\mbox{ otherwise}.$
@@
For $p\geq 2$, the measure $\nu_{p,t}:=\mu_{b_{p,t}}$ is a compactly supported probability measure on $\mathbb{R}$ having for moments
    if $p$ is even,
    $m_{n}(\nu_{p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn/2}{b-1}t^{b},$
    if $p$ is odd,
    $m_{n}(\nu_{p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn/2}{b-1}t^{b}\mbox{ if $n$ even}\quad\quad\mathrm{and}\quad\quad m_{n}(\nu_{p,t})=0\mbox{if $n$ odd},$
@@
    if $p$ is even,
    $\kappa_{n}(\nu_{p,t})=t\mbox{ for all $n\geq 1$},$
    if $p$ is odd,
    $\kappa_{n}(\nu_{p,t})=t\mbox{ for all even $n\geq 1$}\quad\quad\mathrm{and}\quad\quad m_{n}(\nu_{p,t})=0\mbox{ for all odd $n\geq 1$}.$
@@
where $X$ is a real symmetric tensor with entries mean 00, variance $\frac{1}{(p-1)!}$ and bounded moments. This is in particular the case when the entries are Gaussian. On the other hand, a Wishart tensor is a real symmetric tensor of the form
$\mathcal{W}^{k}_{N}:=\frac{x_{1}^{\underline{\otimes}2}+\ldots+x_{k}^{\underline{\otimes}2}}{N^{\frac{p}{2}}},$
where the $x_{i}$ are $(p/2)$-order tensors with i.i.d entries with mean 00, variance $\frac{1}{[(p/2)!]^{2/p}}$ and bounded moments (it is for even $p$, see odd case and formal definitions in Section 3.1). We then have the following convergences.
@@
The last convergence result we state is the free Central Limit Theorem for tensors.
@@
$m_{1}(T)=0\quad\quad\mathrm{and}\quad\quad m_{2}(T)=1\quad\quad\mathrm{and}\quad\quad\forall i,\forall\mathfrak{b}\in\mathcal{M}_{0},|\mathfrak{b}(T_{i})|\leq C(\mathfrak{b}).$
@@
We can also define the $R$-transform which is additive for two freely independent tensors.
Theorem 7.
@@
We point out that if the measures associated to some models of real symmetric tensors converge to compactly supported probability measure, the measure associated to a given tensor has in general a support on $\mathbb{R}$, so we must develop subordination functions and other analytic tools to deal with free convolution in a general setting. It is where the $Q$-transform could be useful. These developments will be the subject of further works, jointly with a better understanding of the link between this measure and the $z$-eigenvalues of a given tensor.
@@ 1.2 Organization of the paper
In Section 2, we will introduce some formalism and prove the analytic moment-cumulant formula. In Section 3, we will define Wigner and Wishart and prove the convergence of their associated laws. In Section 4, we will finally introduce tensorial free convolution for compactly supported measures, $R$-transform and give some basic examples.
@@ 2.1 Maps of tensors
Some examples are given in Figure 1. In particular, we describe some simple maps appearing in the sequel. For $p\geq 1$, the melon maps of degree $p$ (or Frobenius pair in the terminology of [24]), denoted $\mathfrak{f}_{p}^{\sigma}$ for some $\sigma$, are the maps with two vertices and $p$ edges between them : $\pi=(1,\ldots,p)(p+1,\ldots,2p)$ and $\alpha=(1,p+\sigma(1))\cdots(p,p+\sigma(p))$. There are $(p-1)!$ such maps non equivalent, the canonical one being again $\mathfrak{f}_{p}^{\mathrm{id}}$. For $p=2t$ even, the bouquet maps, denoted $\mathfrak{b}_{p}^{\sigma}$ for some $\sigma$, are the ones with a single vertex and $t$ self-loops : $\pi=(1,\ldots,p)$ is a cycle and $\alpha=(\sigma(1),\sigma(2))\cdots(\sigma(p-1),\sigma(p))$. The canonical one is $\mathfrak{b}_{p}^{\mathrm{id}}$. It is a particular case of multicycle map, that is $n\geq 1$ vertices $p/2$ edges between the $i$-th and the $i+1$-th ones for all $1\leq i\leq n$ ($n+1=1$). For a given choice of inputs and outputs edges, there are $[(p/2)!]^{n}$ such maps. For $p$ odd, the definition of an odd multicycle, presented in the introduction follows.
@@ 2.3 Moments of a tensor
We define $\mathcal{B}_{n}$ the set of connected rooted $p$-regular combinatorial maps in $\mathcal{M}_{0}$ with $n$ (unlabeled) vertices.

### v2

@@ 1 Introduction
In many respects, random tensors generalize random matrices to higher orders. This field is currently a vibrant area of research, driven by numerous applications in physics and computer science [2, 23, 10, 4, 22]. In this paper, we aim to establish analogous results within the framework of tensorial freeness. Specifically, we proposed a notion of freeness for tensors in [12], to which we refer for more detailed introduction. Here, we continue developing the foundations of tensorial free probability theory, with a particular focus on the concept of tensorial free convolution. The principal contributions of this work are the following :
@@
where $\gamma$ is the number of connected components of $\mathfrak{b}$. The distribution of a collection of tensors $\mathcal{A}=\{A_{1},\ldots,A_{n}\}$ of possibly different order, $A_{k}\in\mathcal{E}^{N}_{p_{k}}$ is the collection of all trace maps $\mathfrak{b}((T_{v})_{v\in V})$ with $T_{v}\in\mathcal{A}$ and $\mathfrak{b}$ with compatible degrees. In particular, if $T\in\mathcal{E}^{N}_{p}$, the distribution of $T$ is the collection of trace invariants for $p$-regular trace maps, that is all vertices of the map belong to exactly $p$ edges.
We introduce now informally three types of $p$-regular trace maps useful for the rest of the introduction. A melon is a map with two vertices and $p$ edges between them. If $p$ is even, a bouquet is a map with one vertex and $p/2$ self-loops, and a multicycle is a map with $n\geq 1$ vertices $p/2$ edges between the $i$-th and the $i+1$-th ones for all $1\leq i\leq n$ ($n+1=1$). For $p$ odd, a variant of an odd multicycle is presented in Section 2.1.
@@
The notion of freeness detailed in [12] arises from a poset (partially ordered set) on maps, called by analogy the non-crossing poset. Loosely speaking, we say that a map $\mathfrak{b}^{\prime}$ is smaller than a map $\mathfrak{b}$ if they differ by a switch, that is exchanging the ending point of two edges in $\mathfrak{b}$, and $\mathfrak{b}^{\prime}$ has one more connected component. By transitivity, this gives a poset on the set of trace maps with a given number of vertices and a given sequence of degrees. The formal definitions are recalled in Section 2 and an example is given in Figure 2. By Moebius inversion, it is possible to define free cumulants associated to this poset characterized by the implicit relation
@@ 1.1 Main results
Let $\mathcal{B}^{(p)}_{n}$ be the set of connected rooted $p$-regular trace maps with $n$ (unlabeled) vertices. For a tensor $T$ of order $p$, and $n\geq 1$, we define $m_{n}(T)$, the $n$-th moment of $T$, as the sum of $\mathfrak{b}(T,\ldots,T)$ over all $\mathfrak{b}\in\mathcal{B}^{(p)}_{n}$, where all vertices are decorated by $T$. Similarly, $\kappa_{n}(T)$ the $n$-th free cumulant of $T$ is the sum of $\kappa_{\mathfrak{b}}(T,\ldots,T)$ over all $\mathfrak{b}\in\mathcal{B}^{(p)}_{n}$. We set $m_{0}(T)=\kappa_{0}(T)=1$. We then define the corresponding formal power series in $\mathbb{C}{{\left[\left[z\right]\right]}}$ :
@@
We have as functional relation in $\mathbb{C}{{\left[\left[z\right]\right]}}$,
@@
A direct proof of the existence of this measure is still missing.
If $a$ is a distribution on the $p$-regular trace maps (that is $a:\cup_{n}\mathcal{B}^{(p)}_{n}\mapsto\mathbb{R}$ satisfying multilinearity, morphism property and $a(\emptyset)=1$), we may speak of its moments $m_{n}(a)$ as the sum of $a(\mathfrak{b})$ over all $\mathfrak{b}\in\mathcal{B}^{(p)}_{n}$, and we denote $\mu_{a}$ as the measure on $\mathbb{R}$ having $(m_{n}(a))_{n\geq 0}$ as moments, if it exists. By linearity we can equivalently define $a$ by giving $a(\kappa_{\mathfrak{b}})$.
@@
$b_{2p,t}(\kappa_{\mathfrak{b}})=\frac{t}{(p!)^{n}}\mbox{ if }\mathfrak{b}\mbox{ is a multicycle of length }n\mbox{ and }0\mbox{ otherwise,}$,
Note that this is equivalent to $a_{p}(\mathfrak{b})=\mathbf{1}_{\mathfrak{b}\text{ melonic}}$. Moreover, the case of $p$ odd for $b$ is treated in Section 3.1 but not stated here for ease of notation.
@@
$\kappa_{n}(\mu_{p})=1\mbox{ if $n=0$ or $2$}\quad\quad\mathrm{and}\quad\quad\kappa_{n}(\mu_{p})=0\mbox{ otherwise}.$
@@
For $p\geq 1$, the measure $\nu_{2p,t}:=\mu_{b_{2p,t}}$ is a compactly supported probability measure on $\mathbb{R}$ having for moments
$m_{n}(\nu_{2p,t})=\sum_{b=1}^{n}\frac{1}{b}\binom{n-1}{b-1}\binom{pn}{b-1}t^{b},$
@@
$\kappa_{n}(\nu_{2p,t})=t\mbox{ for all $n\geq 1$},$
@@
where $X$ is a real symmetric $p$-order tensor having entries with mean 00, variance $\frac{1}{(p-1)!}$ and bounded moments. This is in particular the case when $X$ belongs to the Gaussian Orthogonal Tensor Ensemble. On the other hand, a $2p$-order Wishart tensor is a real symmetric tensor of the form
$\mathcal{W}^{k}_{N}:=\frac{x_{1}^{\underline{\otimes}2}+\ldots+x_{k}^{\underline{\otimes}2}}{N^{p}},$
where the $x_{i}$ are $p$-order tensors with i.i.d entries with mean 00, variance $\frac{1}{(p!)^{1/p}}$ and bounded moments and $x_{1}^{\underline{\otimes}2}$ is a symmetrization of $x_{i}\otimes x_{i}$. All of that is for even $p$, see the odd case and formal definitions in Section 3.1. We then have the following convergences.
@@
The last convergence result we state now is the free Central Limit Theorem for tensors.
@@
$m_{1}(T_{i})=0\quad\quad\mathrm{and}\quad\quad m_{2}(T_{i})=1\quad\quad\mathrm{and}\quad\quad\forall i,\forall\mathfrak{b}\in\cup_{n}\mathcal{B}_{n},|\mathfrak{b}(T_{i})|\leq C(\mathfrak{b}).$
@@
We may also define the $R$-transform which is additive for two freely independent tensors.
Corollary 3.
@@
We point out that if the measures associated to some models of real symmetric tensors converge to compactly supported probability measure, the measure associated to a given tensor has in general a support on $\mathbb{R}$, so we must develop subordination functions and other analytic tools to deal with free convolution in a general setting. It is where the $Q$-transform could be useful. These developments will be the subject of further works, jointly with a better understanding of the link between this measure and the eigenvalues of a given tensor, which is still unclear.
@@ 1.2 Organization of the paper
In Section 2, we will introduce some formalism and prove the analytic moment-cumulant formula. In Section 3, we will define Wigner and Wishart tensors and prove the convergence of their associated laws. In Section 4, we will finally introduce tensorial free convolution for compactly supported measures, $R$-transform and give some basic examples.
@@ 2.1 Maps of tensors
Some examples are given in Figure 1. In particular, we describe some simple maps appearing in the sequel. For $p\geq 1$, the melon maps of degree $p$ (or Frobenius pair in the terminology of [24]), denoted $\mathfrak{f}_{p}^{\sigma}$ for some $\sigma$, are the maps with two vertices and $p$ edges between them : $\pi=(1,\ldots,p)(p+1,\ldots,2p)$ and $\alpha=(1,p+\sigma(1))\cdots(p,p+\sigma(p))$. There are $(p-1)!$ such maps non equivalent, the canonical one being again $\mathfrak{f}_{p}^{\mathrm{id}}$. For $p=2t$ even, the bouquet maps, denoted $\mathfrak{b}_{p}^{\sigma}$ for some $\sigma$, are the ones with a single vertex and $t$ self-loops : $\pi=(1,\ldots,p)$ is a cycle and $\alpha=(\sigma(1),\sigma(2))\cdots(\sigma(p-1),\sigma(p))$. The canonical one is $\mathfrak{b}_{p}^{\mathrm{id}}$. It is a particular case of multicycle map, that is $n\geq 1$ vertices $p/2$ edges between the $i$-th and the $i+1$-th ones for all $1\leq i\leq n$ ($n+1=1$). For a given choice of inputs and outputs edges, there are $[(p/2)!]^{n}$ such maps. For $p$ odd, an odd multicycle is a map with $2n$ vertices and $(p+1)/2$ edges between the $12i-$ 1-th and the $2i$-th ones, $(p-1)/2$ edges between the $2i$-th and the $12i+$ 1-th ones, for all $1\leq i\leq n$ ($12n+1=$ 1).
@@ 2.3 Moments of a tensor
We remind thet $\mathcal{B}_{n}$ is the set of connected rooted $p$-regular combinatorial maps in $\mathcal{M}_{0}$ with $n$ (unlabeled) vertices.